OpenStudy (anonymous):

Find an equation of the line having the given slope and containing the given point. m=3/4, (7,-9)

OpenStudy (saifoo.khan):

\[y - y_1 = m(x-x_1)\]

OpenStudy (anonymous):

point slope formula right but the answer comes out WEIRD here let me do the problem

OpenStudy (anonymous):

i got y=3/4-30/4

OpenStudy (saifoo.khan):

where'd x go?

OpenStudy (anonymous):

oops y=3/4x-30/4

OpenStudy (saifoo.khan):

checking..

OpenStudy (saifoo.khan):

you should get -57/4 as y-intercept. rest is perfect.

OpenStudy (anonymous):

i dont understand why its -57 please explain

OpenStudy (saifoo.khan):

y - y1 = m(x-x1) m=3/4, (7,-9) y -(-9) = 3/4 ( x - 7)

OpenStudy (anonymous):

I have never seen that version of the formula before. Use: f(x) = mx + b I think that would be easier.

OpenStudy (saifoo.khan):

but this is direct. :P

OpenStudy (anonymous):

wait i still don't understand where the -57 came from

OpenStudy (anonymous):

please tell me where it came from =(

OpenStudy (anonymous):

i did but it came out to be -30/4 i need to understand where and how he got -57 because that's the right answer

OpenStudy (anonymous):

\[(y+9)=3/4(x-7)\] \[(y-9)\times4=3\times(x-7)\] then multiply and you will get the answer

OpenStudy (anonymous):

but thats not part of the formula i am so lost

OpenStudy (anonymous):

y -(-9) = 3/4 ( x - 7) y+9 = 3/4 x - 21/4 y+36/4 = 3/4 x - 21/4 y = 3/4 x - 21/4 -36/4 y = 3/4 x - 57/4

OpenStudy (anonymous):

Do you see it now?

OpenStudy (anonymous):

can you tell me why u multiply 9x4 to get the 36?

OpenStudy (anonymous):

9 = 36/4 You understand that right?

OpenStudy (anonymous):

yes but where di that 36 come from?

OpenStudy (anonymous):

Remember anything over 1 itself 9 = 9/1 I multiplied 4 to the top and bottom side and it doesnt change the values I know that I will subtract 9 from the left side and it will go to the right side where -36/4 There is a rule that in order to add or substract number the denominator has to be the same so 9 has to turn into 36/4

OpenStudy (anonymous):

once 9 turns into 36/4 I can subtract it from - 21/4

OpenStudy (anonymous):

So basically you have to find the lowest common denominator on the left side before you bring it over to the right side to subtract from -21?

OpenStudy (anonymous):

No you can do it after you subtract it from the left side but then I knew I was going to fall into the problem where I had to find the lower common denom. in order to subtract it so I did it beforehand. Sometimes you don't have to do that and the number substract or add much easier.

OpenStudy (anonymous):

These types of problem require you to do a lot of little steps so make sure you write everything out just incase you make a mistake and you can see what you did.

OpenStudy (anonymous):

thank you i get it now

OpenStudy (anonymous):

No problem. Feel free to ask more questions.